OpenStudy (anonymous):

The minute hand of a clock is 5 inches long. What is the area of the circle, in square inches, created as the hand sweeps an hour? is the answer 25?

OpenStudy (anonymous):

i get 6.5 as in 6.5 in^2 per hour

OpenStudy (anonymous):

well the choices are 10, 20, 25, and 100

OpenStudy (anonymous):

or* not and

OpenStudy (mertsj):

The area is 25 pi

OpenStudy (anonymous):

ya... but i didn't think that was the question

OpenStudy (anonymous):

cool, that's what i thought..

hero (hero):

Hint: \[A = \pi r^2\] r = 5 \[A = 25\pi \]

OpenStudy (anonymous):

COOL! that means i was right :) thats good

hero (hero):

I was trying to think of a way to explain without giving the answer, but @mertsj just blurted out the answer

hero (hero):

Actually, I'm just going to report @mertsj

OpenStudy (anonymous):

haha i was just asking if 25 was the right answer

OpenStudy (anonymous):

oh no, don't do that it's fine

OpenStudy (anonymous):

hero is all about you understanding it lol

hero (hero):

Actually it isn't.

hero (hero):

No, I could explain it DoomDude and I was in the process of doing it, but @mertsj blurted out the answer rudely. I don't appreciate it.

OpenStudy (anonymous):

I do not think he meant it maliciously ...

OpenStudy (anonymous):

I already knew the answer....i just wanted someone to tell me if im right

OpenStudy (anonymous):

a confirmation...

OpenStudy (anonymous):

But thx for the help everyone!

hero (hero):

Of course not, he just meant to interrupt and blurt it out. You do realize you're not supposed to post answers only right? @DoomDude @mertsj certainly knows the rules.

OpenStudy (anonymous):

ok, lets just drop this

OpenStudy (anonymous):

this chat is closed